OpenStudy (maeganr):

6 PROBLEMS! HOMEWORK HELP ASAP! QUESTIONS BELOW!

OpenStudy (maeganr):

"Solve each system. Explain why you chose the method you used." #9. 2y=x+1 -2x-y=7 #10. 6x-4y=54 -9x+2y=-69 #11. 3x-2y=8 2x-2y=5 #12. 2x-y=4 3x-y=2 #13. 2x-3y=13 y=(1/2)x-(7/2) #14. -x-3y=-3 2x+3y=5 PLEASE show work ad answer asap! PLEEEEASE.!.

OpenStudy (maeganr):

@erdog82

OpenStudy (anonymous):

@MaeganR Can you please try posting one question per post? It kind of gets overwhelming to the person answering, if you know what i mean :)

OpenStudy (anonymous):

you can use elimination